OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - 3 = x2 + 4x? Answer (-3, 0) and (-1, 0) (3, 0) and (1, 0) (-3, 0) and (1, 0) (3, 0) and (-1, 0)

OpenStudy (mertsj):

Find the x intercepts by letting y = 0 and solve the resulting equation.

OpenStudy (anonymous):

That's where \(y = 0\), so solve\[0-3=x^2+4x \longrightarrow 0=x^2+4x+3\]